# Label Set Optimization via Activation Distribution Kurtosis for Zero-shot Classification with Generative Models

### v1

@@ Abstract
In-context learning (ICL) performance is known to be sensitive to the prompt design, yet the impact of class label options in zero-shot classification has been largely overlooked. This study presents the first comprehensive empirical study investigating how label option (e.g., lexical choice, order, and elaboration) influences zero-shot ICL classification performance. Our findings reveal that lexical choices for label names (e.g., agree vs.support in stance classification) play an important role, with effects also linked to label orders. An analysis of the model internal states further shows that optimal label names tend to activate fewer outlier neurons in the feed forward network. Based on this observation, we propose Label set Optimization via Activation Distribution kurtosiS (LOADS), a post-hoc approach requiring no gradient propagation. LOADS not only demonstrates effectiveness with only 100 unlabelled samples across different model types and sizes, but also shows cross-lingual transferability.
@@ 1 Introduction
Nowadays, generative large language models (LLMs) are widely used directly on classification tasks with zero-shot ICL, where models are prompted to select an option from a pre-defined set of labels (Wang et al., 2022; Antypas et al., 2023; Gonen et al., 2023; Mu et al., 2024). While some classification tasks employ a relatively fixed set of lexicons to represent class labels – such as sentiment analysis (positive and negative) and textual entailment (entailment and contradiction) – other tasks may present more ambiguous choices in lexical selection. Stance classification, for instance, uses diverse pairs of antonyms to represent positive and negative stances across different datasets, e.g. agree-disagree and favor-against. As a result, when crafting prompts for these classification tasks, practitioners face decisions regarding label options in the prompt, such as the lexical selection and the ordering.
@@
Despite studies suggesting the sensitivity of ICL to prompt design (Lu et al., 2022; Yoo et al., 2022; Wei et al., 2024; Mao et al., 2024; Liu et al., 2024a; Zhang et al., 2022; Liu et al., 2022a; Peng et al., 2024; Gonen et al., 2023; Mu et al., 2024), this subtle yet critical consideration of label options in prompt for zero-shot ICL has received limited attention. To fill in this research gap, we explore the impact of three types of label variants (i.e., lexical choice, label order and elaborations) in zero-shot ICL with both encoder-decoder and decoder-only LLMs. We ground our research on stance classification, a task where label adaptation is a known problem due to various label inventories in different studies (Hardalov et al., 2021). We demonstrate that the lexical choice of the label options significantly impacts model performance, also interacting with the label orders. Elaborating on task-related information (e.g. agree with the claim elaborating agree) has minimum effect.
Further analysis of the model internal state is also conducted to gain insights on optimal label selection. Inspired by recent studies on neuron analysis (Kuzmin et al., 2024; Voita et al., 2024; Stolfo et al., 2024), we investigate the neurons in the feed forward network (FFN) in the decoder of the LLMs. We empirically show that prompts with optimal label sets activate fewer outlier neurons. Consequently, we propose a new method, Label set Optimization via Activation Distribution kurtosiS (LOADS), to select optimal label sets for a given classification dataset in zero-shot ICL. LOADS could stably and effectively work with only 100 unlabelled samples of the validation dataset, also demonstrating cross-lingual transferability.
Our contributions can be summarised as follows:
@@
  • Empirical demonstration that zero-shot ICL performance negatively correlates with the number of outlier neurons in FFN when varying the lexical choices for label options. The correlation holds true across diverse English stance classification datasets with different models.
  • A novel and efficient post-hoc method (LOADS) for label selection in zero-shot ICL. Comparing with common strategies in practice, our approach demonstrates statistically significant performance improvements across different models types, model sizes and languages with only 100 unlabeled data samples.
Sections 2, 3 and 4 present our experimental setups, results and discussions for the impact of label options in zero-shot ICL. Then, we describe our neuron analysis of the lexical choice in label options in Section 5 and our proposed method for selecting optimal label sets in Section 6.
@@ 2 Prompting with Varied Label Options for Zero-shot Classification
In zero-shot ICL for classification, a common approach is to provide a set of class label options in the prompt in order to instruct the LLMs to choose one of the options as the classification prediction. Although the label option is a subtle component in the prompt, we are interested in whether it has significant impact on model performance.
Specifically, we explore three types of variants around label options in the prompt: (1) lexical choice; (2) label order; and (3) label elaboration. To properly measure the impact of these factors, we only manipulate with the label options within the same prompt template (Section 3). We show examples of the three variants in Figure 1.
@@
    Expand label sets with WordNet and LLMs. We use WordNet (Fellbaum, 1998) as a reliable source and Claude [^1] https://claude.ai/new as a supplementary source to obtain synonyms for label names in the seed set. For pairs of label names with semantically opposite meanings (such as "agree" and "disagree"), we also consider antonyms to avoid potential ambiguity and present clear contrast for the predicted models.
@@
    Manual selection. We manually select the appropriate label sets generated by Claude to alleviate the impact of noisy label names.
@@
We consider every possible order of the single-word labels in the prompt and compare the model performance against that obtained with the default order. For binary datasets, there is only one alternative arrangement besides the default order, while $N$-way multi-class classification would yield $N!-1$ alternative orders.
@@
We investigate whether transforming single-word labels (e.g., "agree") into more detailed phrases (e.g., "agree with the claim") has an impact on model performance. On the one hand, elaborating on task details with the label may provide the model with a stronger alignment signal between the label and the task, emphasizing what the label is referring to. On the other hand, it also increases the label length and may introduce noise in the prompt (Liu et al., 2024a). Therefore, we design three levels of elaborations (shorted for E1, E2 and E3) by progressively adding more task-related (and potentially redundant) information to the single-word label names, as shown in Figure 1.
3 Experimental Setups
@@
We focus on the stance classification task due to its rich label inventories across various readily available datasets. We consider four binary and five multi-class datasets from existing English stance classification benchmarks (Schiller et al., 2021; Hardalov et al., 2021; Chen et al., 2023), as shown in Table 1. The nine datasets cover different domains, such as social media posts, news articles and online debates forums. The target of the stance varies from non-phrase topics (e.g., "immigration" or "school uniform") to phrases or statements (e.g., news headlines, claims or rumourous tweets). Details can be found in Appendix A.
Dataset Name Labels
scd
(Hasan and Ng, 2013)
original: for, against
best: pro, con
perspectrum
(Chen et al., 2019)
original: support, undermine
best: validate, refute
snopes
(Hanselowski et al., 2019)
original: agree, refute
best: affirm, refute
ibmcs
(Bar-Haim et al., 2017)
original: pro, con
best: endorse, deny
vast
(Allaway and McKeown, 2020)
original: pro, con, neutral
best: confirm, dispute, neither
emergent
(Ferreira and Vlachos, 2016)
original: for, against, observing
best: endorse, reject, neutral
semeval
(Mohammad et al., 2016)
original: favour, against, neither
best: accept, reject, neutral
rumoureval
(Gorrell et al., 2019)
original: support, deny, query, comment
best: confirm, reject, question, neutral
arc
(Hanselowski et al., 2018)
original: agree, disagree, discuss, unrelated
best: affirm, refute, discuss, unrelated
Table 1: Lists of the English stance classification datasets, their labels in original dataset, and the best labels on Flan-T5-xl with zero-shot ICL as an example to justify our motivation on LOADS.
@@
We cover both encoder-decoder and decoder-only LLMs, and experiment with the prevalent open-sourced Flan-T5 (Chung et al., 2024), Llama 3 and Llama 3.1 (Dubey et al., 2024) model families as representatives for these two types of LLMs. We choose their moderate-sized instruction-tuned versions, Flan-T5-xl (3b), Llama -3-Instruct (8b) and Llama-3.1-Instruct (8b), as our primary models for investigation due to our hardware resources constraints and their decent zero-shot ICL performance (Aiyappa et al., 2024; Chung et al., 2024; Dubey et al., 2024).
@@
Table 2: The maximum (max), minimum (min), average (avg), variance (var) of the model performance across different label names in the prompt for each validation set. The performance of the original label set (Original) is also included, showing that they fail to reach the maximum performance LLMs could get. The extent of the gap between the maximum and minimum performances is represented using colors: max--min$>0.3$, $0.2<$max--min$<0.3$. The greater the variance, the greater the impact of label lexical choice, and the greater the potential utility of optimizing the label set.
@@
We refer to the prompt template used in the supervised fine-tuning of Flan-T5 and Llama (Wang et al., 2022; Chung et al., 2024; Dubey et al., 2024). We fine-tune the template on validation sets to ensure the LLMs could follow the instructions on more than 95% of the samples. We also test other templates and find the results are consistent (e.g., prompting with label explanations in Appendix E). We present the results with the following prompt template in this paper: Given a [text1_name] and a [text2_name], detect the stance that the [text2_name] has towards the [text1_name]. There are {N} options: "{label0}", "{label1}, … , and {labelN-1}". Now complete the following example. [text1_name]: {text1}. [text2_name]: {text2}.
@@
[…]
@@
To ensure reproducibility, we use greedy search for decoding [^3] We present the potential impact of the decoding strategy in Appendix D.. In more than 95% cases, LLMs exactly follow the instruction and output the stance name within the required stance options. Therefore, we directly use the model generation as the predicted label without post-processing or mapping. We run experiments on the validation set of each dataset. See Appendix C for details on the label sets experimented with. We exclude the topic/entities-based (such as stance towards Obama) stance classification datasets (i.e. scd, semeval and vast) in label elaboration experiments to avoid unnecessary ambiguity or bias during elaboration (e.g., the text to be classified may refer to anything from policy to personal behavior about Obama, and elaborating agree to agree with the opinion of Obama could lead to biased prediction).
4 Results and Analysis
We first discuss the impact of lexical choice, label order, and label elaboration on zero-shot ICL for classification (Subsection 4.1). Then we analyze whether optimal or poor label options could be generalized across datasets and models (Subsection 4.2). Suggestions for practitioners in zero-shot ICL for classification appear in Subsection 4.3.
4.1 Results
(a) top-k optimal label sets
(b) top-k poor label sets
Figure 2: The maximum performance gain (positive value) and drop (negative value) on each validation set after re-ordering the label names for the top-k optimal (a) and poor (b) label sets with Llama 3. Results on other two LLMs can be found in Appendix F.
@@ 4.1 Results
Overall, we observe limited influence of label order (see Appendix F). However, we are also interested in extremes of performance change caused by shifting label orders, particularly the maximum performance gain and drop, and whether the extreme gain or drop correlates with the lexicons used for the label names. Therefore, for each dataset, we first select the top-k [^4] Due to the exponential increasing of label order options and our limited computational resources, we set k=15 for binary classification; k=10 for three-way classification and k=2 for four-way classification optimal and poor label sets based on their performance in the same order (i.e. the default order). We then examine the maximum increase or decrease of the performance after re-arranging the label orders for optimal and poor label sets, respectively. As shown in Figure 2(a), altering the order for the optimal label sets has the risk of high performance drops (e.g., even more than 0.2 on the binary classification snopes dataset). While performance improvements are possible, the gains are relatively limited (lower than 0.1 on all datasets). Conversely, re-arranging the order for the poorly performing label sets offers potential for substantial improvement (Figure 2(b)). This high improvements for certain label sets after re-ordering suggests that the poor performance may partly stem from the initial sub-optimal ordering.
@@
Similarly, we select the top-k555k=15 for binary classification, k=30 for three/four-way classification. optimal and poor single-word label sets based on their performance, and examine the performance change after elaborations. As shown in Table 3, the models are robust to the elaborations for either optimal or poor single-name label sets, indicating that adding task related details or increasing the label token lengths brings limited impact on average. However, we also observe relatively large performance change on certain datasets. Specifically, for the rumoureval dataset with Llama 3, we notice a performance drop larger than 0.2 when elaborating an optimal label set. Performance on the ibmcs dataset with Llama 3.1 could increase 0.2 when elaborating a poorly-performing label set.
Dataset $E_{1}$ $E_{2}$ $E_{3}$
Opt. Sub-opt. Opt. Sub-opt. Opt. Sub-opt.
perspectrum 0.016 0.018 0.010 0.015 0.017 0.009
ibmcs 0.027 0.041 0.024 0.014 0.029 0.044
snopes 0.055 0.040 0.054 0.026 0.029 0.018
emergent 0.051 0.053 0.047 0.038 0.022 0.040
rumoureval 0.095 0.036 0.084 0.020 0.058 0.102
arc 0.017 0.024 0.015 0.021 0.035 0.049
Table 3: The average absolute performance change after elaborating for optimal or poor single-word label sets with Llama 3 ($E_{1}$, $E_{2}$, $E_{3}$ see Figure 1). Results of other LLMs could be found in Appendix G.
4.2 Analysis
A natural question arising from the observed substantial impact of the lexical choices for label options in zero-shot ICL is: Can optimal or poor label names be generalised across datasets or even models? If so, for any new datasets/models, we could potentially adopt the optimal label names we found on related dataset/models. Therefore, for each dataset $D_{i}$ on each LLM $M_{j}$, we select the top-$k$ label sets with optimal ($L^{optimal}_{D_{j}\_M_{j}}$) or poor ($L^{poor}_{D_{i}\_M_{j}}$) performance, and calculate the number of overlap between the label sets. We only focus on the positive and negative stances in this section to enable comparison across binary, three-way and four-way stance classification datasets.
Generalisation across datasets within the same model.
To analyse whether the optimal/poor label sets found on one dataset $D_{i}$ could be adapted to another similar dataset $D_{j}$ on the same LLM $M$, we calculate the overlap of top-k optimal/poor ($L_{D_{i}\_M}\bigcap L_{D_{j}\_M}$) between each dataset.
Our analysis reveals that there is no guarantee of generalizability for optimal/poor label sets across stance datasets on the same model. The level and type of generalization across datasets are also model-dependent. For Llama 3.1, there is always an overlap among the worst top-5 label sets across all the datasets, but not among the best top-5. That is, for all datasets tested, there are shared label sets that always result in sub-optimal performance when using Llama 3.1. Also, using supportive and opposed as label options leads to the worst performance on seven out of nine datasets – Llama 3.1 tends to overly generate supportive. However, no single label set consistently yields the best performance across more than half of the datasets with Llama 3.1. Flan-T5 follows a similar trend.
Conversally, for Llama 3, the level of transferability of the optimal and poor label sets across datasets is relatively similar to each other. Certain datasets, e.g. semeval, exhibit low transferability of optimal and poor label sets from other datasets (it overlaps with only four/three datasets considering top-5 label sets), while for vast, its top-5 optimal or poor label sets overlap with every dataset we tested when using Llama 3. Overall, the level of transferability for optimal label sets on Llama 3 is higher than Llama 3.1 and Flan-T5.
Generalisation across models with the same dataset.
To analyse whether the optimal/poor label sets found for dataset $D$ on model $M_{i}$ could be adapted to another model $M_{j}$, we calculate the overlap of top-k optimal/poor ($L_{D\_M_{i}}\bigcap L_{D\_M_{j}}$) between each model.
We find that optimal or poorly-performing label sets identified for Llama 3 on a specific dataset are more likely to transfer effectively to Llama 3.1 and Flan-T5, compared to transfer from Llama 3.1 to Flan-T5. However, the degree of generalisation varies depending on the dataset. We include the full statistics in Appendix H.
perspectrum ibmcs snopes scd emergent semeval rumoureval arc vast
Llama 3 (8b) -0.4921* -0.3787* -0.4359* -0.4217* -0.5781 -0.2618* -0.3764* -0.1662 -0.3639*
Llama 3.1 (8b) -0.4103* -0.3642* -0.1874 -0.0686 -0.4310* -0.2232* -0.3944* -0.1708* -0.1208
Flan-T5-xl (3b) -0.4476* -0.3638* -0.6014* 0.2353 -0.1089 -0.2881* -0.1714* -0.5742 0.1003

### v2

@@ Abstract
In-context learning (ICL) performance is highly sensitive to prompt design, yet the impact of class label options (e.g. lexicon or order) in zero-shot classification remains underexplored. This study proposes LOADS (Label set Optimization via Activation Distribution kurtosiS), a post-hoc method for selecting optimal label sets in zero-shot ICL with large language models (LLMs). LOADS is built upon the observations in our empirical analysis, the first to systematically examine how label option design (i.e., lexical choice, order, and elaboration) impacts classification performance. This analysis shows that the lexical choice of the labels in the prompt (such as agree vs. support in stance classification) plays an important role in both model performance and model’s sensitivity to the label order. A further investigation demonstrates that optimal label words tend to activate fewer outlier neurons in LLMs’ feed-forward networks. LOADS then leverages kurtosis to measure the neuron activation distribution for label selection, requiring only a single forward pass without gradient propagation or labelled data. The LOADS-selected label words consistently demonstrate effectiveness for zero-shot ICL across classification tasks, datasets, models and languages, achieving maximum performance gain from 0.54 to 0.76 compared to the conventional approach of using original dataset label words.
@@ 1 Introduction
Generative large language models (LLMs) are increasingly used for classification tasks via zero-shot in-context learning (ICL), where models are prompted to select an option from a pre-defined set of labels Wang et al. (2022); Antypas et al. (2023); Gonen et al. (2023); Mu et al. (2024). While some classification tasks employ a relatively fixed set of lexicons to represent class labels, such as sentiment analysis (positive and negative) and textual entailment (entailment and contradiction), other tasks may present more ambiguous choices in lexical selection. Stance classification, for instance, uses diverse pairs of antonyms to represent positive and negative stances across different datasets, e.g. agree-disagree vs. favor-against. As a result, when crafting prompts for these classification tasks, practitioners face decisions regarding label options in the prompt, such as lexical selection and ordering.
@@
Despite studies suggesting the sensitivity of ICL to prompt design Lu et al. (2022); Yoo et al. (2022); Wei et al. (2024); Mao et al. (2024); Liu et al. (2024a); Zhang et al. (2022); Liu et al. (2022a); Peng et al. (2024); Gonen et al. (2023); Mu et al. (2024), this subtle yet critical consideration of label options in prompt for zero-shot ICL has received limited attention. To fill in this research gap, we explore the impact of three types of label variants (i.e., lexical choice, label order and elaborations) in zero-shot ICL with both encoder-decoder and decoder-only LLMs. We mainly ground our research on stance classification, a task where label adaptation is a known problem due to various label inventories in different studies Hardalov et al. (2021). We demonstrate that the lexical choice of the label options significantly impacts model performance. The model’s sensitivity to the label order also depends on the lexical choice, while elaborating on task-related information (e.g. agree with the claim elaborating agree) has minimum effect.
Inspired by recent studies on neuron analysis Kuzmin et al. (2023); Voita et al. (2024); Stolfo et al. (2024), we further investigate the neurons in the feed-forward network (FFN) in the decoder of the LLMs. We empirically show that prompts with optimal label sets activate fewer outlier neurons. Consequently, we propose a new method, Label set Optimization via Activation Distribution kurtosiS (LOADS), to select optimal label sets for a given classification dataset in zero-shot ICL. LOADS could stably and effectively work with only 100 unlabelled samples of the validation dataset, also demonstrating transferability across datasets and languages. Our contributions are summarized as follows:
@@
  • The empirical demonstration that zero-shot ICL performance negatively correlates with the number of outlier neurons in FFN when varying the lexical choices for label options. The correlation holds true across diverse English stance classification datasets and topic classification datasets with different models.
  • A novel and efficient post-hoc method (LOADS) for label selection in zero-shot ICL. Compared with common strategies in practice, our approach demonstrates statistically significant performance improvements across model types, model sizes and languages with only 100 unlabeled data samples. Our analysis also suggests that the LOADS-selected label set is potentially transferable across similar datasets for a specific LLM, further alleviating the cost to collect samples for the target new dataset.
We present our experimental setups, results and discussions on the impact of label options in zero-shot ICL in Section 3. Then, we describe our neuron analysis of the lexical choice in label options in Section 4 and our proposed method LOADS for selecting optimal label sets in Section 5.
2 Related Work
ICL Performance
Few-shot ICL mainly focuses on cases where LLMs are directly prompted with $N$ demonstration examples. The studies highlight the substantial impact of example ordering Lu et al. (2022), formatting Yoo et al. (2022); Wei et al. (2024); Mao et al. (2024); Liu et al. (2024a), and examples selection Zhang et al. (2022); Liu et al. (2022a); Peng et al. (2024). The parallel lines of work focus on improving few-shot ICL via the optimal selection or arrangement of examples Liu et al. (2022a); Rubin et al. (2022); Lu et al. (2022); Zhang et al. (2022); Liu et al. (2024b); Xu et al. (2024), re-weighting examples Yang et al. (2023), automatic reformat or generation of demonstration representations Kim et al. (2022); Liu et al. (2023), and introduction of intermediate reasoning steps Wei et al. (2022); Zhang et al. (2023b). However, the impact of lexical choices for label names in classification received little attention, with the only closely related work suggesting that LLMs are likely to confuse classes which share similar key vectors in the attention modules Wang et al. (2023).
For zero-shot ICL, Mu et al., Mu et al. (2024) demonstrate the effect of using synonyms for class options, but they neither consider the order of the label options nor propose an effective strategy to choose the label names. Gonen et al., Gonen et al. (2023) empirically show that perplexity could be an effective indicator for prompt selection, but they do not account for class options. Notably, behavioral differences between few-shot and zero-shot ICL have been frequently observed, suggesting that findings from few-shot ICL do not necessarily hold in the zero-shot context Lin and Lee (2024).
Prompt-Tuning and Verbalizer
Prompt-tuning aims to automatically find or generate an optimal discrete prompt (e.g., through gradient-based search Shin et al. (2020); Shi et al. (2023) or fine-tuning Gao et al. (2021); Le Scao and Rush (2021); Deng et al. (2022)) or by training continuous prompt Lester et al. (2021); Liu et al. (2022b). Verbalizer can be taken as a mapping function that links discrete class labels to corresponding tokens or phrases in a model’s vocabulary. A range of methods developed to build the verbalizer, including manually created verbalizer Schick and Schütze (2021), search-based verbalizer that identifies label words automatically from the dataset Gao et al. (2021); Shin et al. (2020), and soft verbalizers that uses continuous embeddings obtained through fine-tuning Hambardzumyan et al. (2021); Cui et al. (2022). Prompt-tuning often does not focus on label set selection for zero-shot ICL, and the verbalizer introduces additional components to the decoding of the generative models, distinguishing it fundamentally from our work.
@@ 3 Prompting with Varied Label Options for Zero-shot Classification
In zero-shot ICL for classification, a common approach is to provide a set of class label options in the prompt to instruct the LLMs to choose one of the options as the classification prediction. Although the label option is a subtle component in the prompt, we are interested in whether it has a significant impact on model performance.
Specifically, we explore three types of variants around label options in the prompt: (1) lexical choice; (2) label order; and (3) label elaboration. To accurately measure the impact of these factors, we only manipulate the label options within the same prompt template (Section 3.2). We show examples of the three variants in Figure 1.
3.1 Methodology
@@
    Expand label sets with WordNet and LLMs. We use WordNet Fellbaum (1998) as a reliable source and Claude [^1] https://claude.ai/new as a supplementary source to obtain synonyms for label names in the seed set. For pairs of label names with semantically opposite meanings (such as “agree" and “disagree"), we also consider antonyms to avoid potential ambiguity and present clear contrast for the predicted models.
@@
    Manual selection. We manually filter out semantically unrelated or inappropriate label sets generated by Claude to mitigate the impact of noisy label names.
@@
We consider every possible order of the single-word labels in the prompt and compare the model performance against that obtained with the default order. For binary datasets, there is only one alternative arrangement besides the default order, while $N$-way multi-class classification would yield N!1N!-1 alternative orders.
@@
We investigate whether transforming single-word labels (e.g., “agree") into more detailed phrases (e.g., “agree with the claim") has an impact on model performance. On the one hand, elaborating on task details with the label may provide the model with a stronger alignment signal between the label and the task, emphasizing what the label is referring to. On the other hand, it also increases the label length and may introduce noise in the prompt Liu et al. (2024a). Therefore, we design three levels of elaborations (shorted for E1, E2 and E3) by progressively adding more task-related (and potentially redundant) information to the single-word label names, as shown in Figure 1.
3.2 Experimental Setups
@@
We focus on the stance classification task due to its rich label inventories across various readily available datasets. Stance classification aims to identify the type of an expressed opinion (e.g., “agree" or “disagree") in a given piece of text towards a particular topic, claim, or entity. We consider four binary (scd Hasan and Ng (2013), perspectrum Chen et al. (2019), snopes Hanselowski et al. (2019) and ibmcs Bar-Haim et al. (2017)) and five multi-class datasets (vast Allaway and McKeown (2020), emergent Ferreira and Vlachos (2016), semeval Mohammad et al. (2016), rumoureval Gorrell et al. (2019) and arc Hanselowski et al. (2018)) from existing English stance classification benchmarks Schiller et al. (2021); Hardalov et al. (2021); Chen et al. (2023), as shown in Table 1. The nine datasets cover different domains, such as social media posts, news articles and online debates forums. We also experiment with topic classification in Section 4 and 5 to demonstrate the generalizability of our findings to other NLP tasks.
Dataset Name Original Label Words Optimal Label Words
scd for, against pro, con
perspectrum support, undermine validate, refute
snopes agree, refute affirm, refute
ibmcs pro, con endorse, deny
vast pro, con, neutral confirm, dispute, neither
emergent for, against, observing endorse, reject, neutral
semeval favour, against, neither accept, reject, neutral
rumoureval support, deny, query, comment confirm, reject, question, neutral
arc agree, disagree, discuss, unrelated affirm, refute, discuss, unrelated
Table 1: Lists of the English stance classification datasets, their labels in original dataset, and the optimal labels with the highest zero-shot ICL performance on Flan-T5-xl as an example to justify our motivation on LOADS.
@@
We cover both encoder-decoder and decoder-only LLMs, and experiment with the prevalent open-sourced Flan-T5 Chung et al. (2024), Llama 3 and Llama 3.1 Dubey et al. (2024) model families as representatives for these two types of LLMs. We choose their moderate-sized instruction-tuned versions, Flan-T5-xl (3b), Llama -3-Instruct (8b) and Llama-3.1-Instruct (8b), as our primary models for investigation due to our hardware resources constraints and their decent zero-shot ICL performance Aiyappa et al. (2024); Chung et al. (2024); Dubey et al. (2024). We conduct experiments with Gemma-2-it (9b)  [^2] https://huggingface.co/google/gemma-2-9b-it and Flan-T5-xxl (13b) to further show the generalizability of LOADS in Section 5.
@@
Table 2: The maximum (max), minimum (min), average (avg), variance (var) of the model performance across different label names in the prompt for each validation set. The performance of the original label set (Original) is also included, showing that they fail to reach the maximum performance LLMs could get. The extent of the gap between the maximum and minimum performances is represented using colors: max-min>0.3>0.3, 0.2<0.2<max-min<0.3<0.3. The greater the variance, the greater the impact of label lexical choice, and the greater the potential utility of optimizing the label set.
@@
We refer to the prompt template used in the supervised fine-tuning of Flan-T5 and Llama Wang et al. (2022); Chung et al. (2024); Dubey et al. (2024). We present the results with the following prompt template in this paper: Given a [text1_name] and a [text2_name], detect the stance that the [text2_name] has towards the [text1_name]. There are {N} options: “{label0}", “{label1}, … , and {labelN-1}". Now complete the following example. [text1_name]: {text1}. [text2_name]: {text2}. We also test other templates and find the results are consistent (e.g., prompting with label explanations in Appendix E).
@@
We adopt macro-$F1$ for model performance evaluation to align with prior studies Schiller et al. (2021); Hardalov et al. (2021); Chen et al. (2023). We use wF2wF2333wF2wF2 gives different weights for each stance: deny == support =0.40=0.40, query =0.15=0.15 and comment =0.05=0.05. to account for data imbalance in rumoureval Scarton et al. (2020).
@@
To ensure reproducibility, we use greedy search for decoding [^4] Potential impact of the decoding strategy can be found in the Appendix D.. In more than 95% cases, LLMs exactly follow the instruction and output the stance name within the required stance options. Therefore, we directly use the model generation as the predicted label without post-processing or mapping. We run experiments on the validation set of each dataset. See Appendix C for details on the label sets experimented with. We exclude the topic/entities-based (such as stance towards Obama) stance classification datasets (i.e. scd, semeval and vast) in label elaboration experiments to avoid unnecessary ambiguity or bias during elaboration (e.g., the text to be classified may refer to anything from policy to personal behavior about Obama, and elaborating agree to agree with the opinion of Obama could lead to biased prediction).
3.3 Results and Analysis
We first discuss the impact of lexical choice, label order, and label elaboration on zero-shot ICL for classification. Then we provide suggestions for practitioners in zero-shot ICL for classification.
@@ 3.3 Results and Analysis
On average, we observe limited influence of label order (see the Appendix F for details). However, we are also interested in extremes of performance change caused by shifting label orders, particularly the maximum performance gain and drop, and whether the extreme gain or drop correlates with the lexicons used for the label names. Therefore, for each dataset, we first select the top-k [^5] Due to the exponential increasing of label order options and our limited computational resources, we set k=15 for binary classification; k=10 for three-way classification and k=2 for four-way classification optimal and poor label sets based on their performance in the same order (i.e. the default order). We then examine the maximum increase or decrease of the performance after re-arranging the label orders for optimal and poor label sets, respectively.
We find that altering the order for the optimal label sets has the risk of high performance drops (e.g., even more than 0.2 on the binary classification snopes dataset, Figure 2(a) in Appendix). While performance improvements are possible, the gains are relatively limited (lower than 0.1 on all datasets). Conversely, re-arranging the order for the poorly performing label sets offers potential for substantial improvement (Figure 2(b)). This high improvements for certain label sets after re-ordering suggests that the poor performance may partly stem from the initial sub-optimal ordering.
@@
Similarly, we select the top-k666k=15 for binary classification, k=30 for three/four-way classification. optimal and sub-optimal single-word label sets based on their performance, and examine the performance change after elaborations. We observe that the models are robust to the elaborations for either optimal or poor single-name label sets (full results in Table 9), indicating that adding task related details or increasing the label token lengths brings limited impact on average. However, we also observe relatively large performance change on certain datasets. Specifically, for the rumoureval dataset with Llama 3, we notice a performance drop larger than 0.2 when elaborating an optimal label set. Performance on the ibmcs dataset with Llama 3.1 could increase 0.2 when elaborating a poorly-performing label set.
Model Stance Classificaion Topic Classification
perspectrum ibmcs snopes scd emergent semeval rumoureval arc vast TweetTopic AG News
Llama 3 (8b) -0.4921* -0.3787* -0.4359* -0.4217* -0.5781 -0.2618* -0.3764* -0.1662 -0.3639* -0.4994* -0.2447*
Llama 3.1 (8b) -0.4103* -0.3642* -0.1874 -0.0686 -0.4310* -0.2232* -0.3944* -0.1708* -0.1208 -0.2196* 0.0200
Flan-T5-xl (3b) -0.4476* -0.3638* -0.6014* 0.2353 -0.1089 -0.2881* -0.1714* -0.5742 0.1003 0.1587 0.0398